# Generalized Petermann factor of non-Hermitian systems at exceptional points

### v1

@@ Abstract
The nonorthogonality of modes in open systems significantly modifies their resonant response, resulting in quantitative and qualitative deviations from Breit-Wigner resonance relations. For isolated resonances with a Lorentzian lineshape, the deviations amount to a quantitative enhancement of the resonance linewidth, given by the Petermann factor, which is determined by the overlap of left and right eigenmodes of the underlying effectively non-Hermitian Hamiltonian. The Petermann factor diverges at exceptional points (EP), where complex resonance frequencies become degenerate, and right and left eigenmodes are orthogonal to each other. This divergence signifies a qualitative departure from a Lorentzian lineshape, which has moved into the focus of recent attention. In this work, we develop the analogue of the Petermann factor for this qualitatively modified response at an EP, and describe how this EP Petermann factor manifests in a variety of physical scenarios. Firstly, we identify this analogue in physical terms as an enhancement of the response of a system to external or parametric perturbations. Utilizing two natural orthogonally projected reference systems based on the right and left eigenvectors, we show that each choice carries a precise geometric interpretation that naturally extends the notion of the Petermann factor for isolated resonances to EPs. The two choices can be combined into an overall EP Petermann factor, which again can be expressed in purely geometric terms. Secondly, we illuminate the geometric mechanisms that determine the size of the EP Petermann factor, by considering the role of modes participating in the degeneracy and those that remain spectrally separated from it. This also leads to a systematic description to evaluate this factor. Thirdly, we design a system that allows us to study the EP Petermann factor in a specific physical setup, consisting of two microrings coupled to a waveguide with embedded semitransparent mirrors. With this example, we demonstrate that our approach gives more accurate results for the spectral response strength than conventional truncation schemes. These results complete the description of systems operating at exceptional points in the same way as the original Petermann factor does for isolated resonances, and pave the way to the design of open systems with unconventional spectral response, be it in emission or in response to static and dynamic perturbations.
@@ I Introduction
The field of non-Hermitian physics has attracted a considerable amount of attention in various subdisciplines, such as photonics [1, 2], acoustics [3], and condensed-matter physics [4]. The term “non-Hermitian” refers to the (effective) Hamiltonian [5, 6], which does not only describe conservative but also dissipative dynamics of waves, stemming, e.g., from material gain, material loss, and radiation loss.
@@
The Petermann factor diverges to infinity [24, 25, 26] at so-called exceptional points (EPs) in parameter space [27]. At such an EP of order $n$ exactly $n$ eigenstates and their corresponding eigenvalues of the Hamiltonian coalesce. EPs are highly sensitive to perturbations: a perturbation of strength $\varepsilon>0$ results generically in eigenvalue changes proportional to the $n$th root of $\varepsilon$ [27]. This sensitivity makes EPs a good candidate for a new type of sensors [28, 29, 30, 31, 32, 33, 34, 35]. The response of the system at the EP in terms of eigenvalue changes can be quantified by the spectral response strength $\xi$ [36].
The aim of this paper is to extend the concept of the Petermann factor, which was originally introduced for isolated resonances, to the degenerate situation at an EP. Physically, this EP Petermann factor quantifies the enhancement of the response strength above that of reference systems from which the mode-nonorthogonality is removed. For these systems there are two natural choices: one pertains to the excitation of the system, while the other relates to the detection of the resulting response. The corresponding enhancement factors carry a precise geometric interpretation that naturally extends the notion of the Petermann factor for isolated resonances to EPs, and link the physical enhancement of the response strength to purely geometric information quantifying nonorthogonality of the right and left generalized eigenvectors. This information can further be combined into an overall EP Petermann factor, which again carries a precise geometric interpretation quantifying mode non-orthogonality. To substantiate the physical interpretation of these concepts, we illuminate the role of the EP Petermann factor in various contexts, such as system response to perturbations and quantum noise. In addition to these theoretical insights, we present a detailed prescription for determining the EP Petermann factor from both analytical and numerical models. Finally, we demonstrate the practical importance of the EP Petermann factor by comparing its predictions for the spectral response strength to that of standard truncation schemes. Through both an analytical toy model and a full numerical simulation of a realistic photonic system, we reveal how the inclusion of the EP Petermann factor leads to significantly more accurate results.
@@ II.1 Definitions and preliminary considerations
In this work, we consider the consequences of mode nonorthogonality for the physical response of systems described by effectively non-Hermitian Hamiltonians $\hat{H}$ in different spectral scenarios. We do not make any specific assumptions about the nature of the effective Hamiltonian, i.e., whether it arises from a mean-field description, post selection, or elimination of unobserved parts or freedoms, but assume for concreteness that the relevant part of the point spectrum can be described by an $N\times N$-dimensional matrix $H$. In a given orthonormal basis $\{|n\rangle\}_{1\leq n\leq N}$, $\langle n|m\rangle=\delta_{nm}$, the elements of this matrix are given by $H_{nm}=\langle n|\hat{H}|m\rangle$. We denote the identity operator as $\hat{I}$, and the $N\times N$-dimensional identity matrix as $𝟙\openone$.
@@ II.2 Spectral decomposition of Hermitian matrices
We recall that for a Hermitian Hamiltonian, the eigenvalues $E_{l}$ are real, while the right and left eigenvector can be identified via $\langle L_{l}|=|R_{l}\rangle^{\dagger}$. Additionally, the eigenvectors of distinct eigenvalues $E_{k}\neq E_{l}$ are mutually orthogonal, $\langle R_{k}|R_{l}\rangle=0$, and individual eigenvectors can be normalized to obey $\langle R_{l}|R_{l}\rangle=1$. Furthermore, this construction can be straightforwardly extended to the case of degenerate eigenvalues. The mathematical reason is that in Hermitian systems, the algebraic multiplicity $n_{l}$, given by the multiplicity of roots in the secular equation (2), is identical to the geometric multiplicity $g_{l}$, given by the number of linearly independent eigenvectors admitted by Eq. (1). Employing a suitable orthogonal basis in the subspace spanned by these vectors, we therefore can construct an orthonormal basis for the complete systems, with
[…]
@@
[…]
@@
[…]
@@
The spectra decomposition then takes the form
[…]
@@ II.4 Exceptional points
Degeneracies of non-Hermitian systems differ from degeneracies in Hermitian systems in that along with the eigenvalues, the eigenvectors of the participating eigenvalues can also coalesce. For typical degeneracies in non-Hermitian systems, the geometric multiplicity is indeed smaller than the algebraic multiplicity, and in the generic case (i.e., without further fine tuning of parameters), the geometric multiplicity is one. These generic non-Hermitian degeneracies are called exceptional points (EPs). To cover these cases, we can view the indices $l$ as labels of separate eigenvectors, each associated with $n_{l}$ degenerate eigenvalues. This includes the case that the spectrum displays several EPs, while scenarios with higher geometric multiplicity can, in the context of the present work, be considered as additional accidental degeneracies between different EPs.
As the number of independent eigenvectors is reduced, these vectors do not form a complete basis (the Hamiltonian is then said to be defective). An important consequence is that the Hamiltonian can no longer be diagonalized. Instead, a complete basis can obtained by supplementing each right eigenvector $|R_{l}\rangle$ by a Jordan chain of $n_{l}-1$ covectors $|J^{(l)}_{n}\rangle$, fulfilling
@@
where nominally $|J^{(l)}_{1}\rangle=|R_{l}\rangle$. Analogously, one can define a chain of left covectors $\langle\tilde{J}^{(l)}_{m}|$, which fulfill
@@
The transformation into this biorthogonal basis corresponds to a Jordan decomposition
@@
where $D$ is the matrix with unit entries on the first superdiagonal, i.e., $D_{nn^{\prime}}=\delta_{n,n^{\prime}-1}$. An important feature of these matrices are that they are nilpotent with index $n_{l}$, hence, $D^{n_{l}}=0$ but $D^{n_{l}-1}\neq 0$. For an isolated eigenvalue, we simply have $D=0$, so that this term can be dropped from the decomposition, in agreement with Eq. (10).
@@
While the construction of the biorthogonal Jordan chains allows for some freedom, which is also reflected in the detailed form of the transformation matrix $T$, the operators $\hat{P}_{l}$ and $\hat{N}_{l}$ are unique. From their construction, the operators $\hat{P}_{l}$ are oblique projectors that continue to fulfill Eq. (6), but now are of finite rank $n_{l}$. Furthermore, the operators $\hat{N}_{l}$ are nilpotent with index $n_{l}$, i.e., $\hat{N}_{l}^{n_{l}}=0$ but $\hat{N}_{l}^{n_{l}-1}\neq 0$. Additionally, these operators obey the important relation
@@ II.5 Measures of mode nonorthogonality for isolated resonances
The spectral decompositions described above imply a strong link between the spectral features of non-Hermitian systems and the geometric structure of the states associated with them. To prepare the discussion of these features at exceptional points, we survey common theoretical mode-nonorthogonality measures that illuminate different aspects of this link for spectrally isolated resonances, and detail how they are closely related to each other. The detailed physical contexts for these quantities are provided in later sections, where we generalize these notions to apply to systems at exceptional points.
@@
[…]
@@
[…]
@@
Even though formulated from different perspectives, the different mode-nonorthogoality measures are therefore strictly related to each other. Hence, for isolated resonance, physical considerations of mode crosstalk, wavefunction imbalances, and eigenvalue sensitivity all relate to a very specific and universal aspects of mode nonorthogonality. In the following sections, we will explore how a richer and more complex picture emerges at EPs.
@@ II.6 Mode nonorthogonality near exceptional points
$\xi_{l}=||\hat{N}_{l}^{n_{l}-1}||_{2}$ (25)
arises from the nilpotent part of the spectral decomposition at the EP, with the spectral norm $||\hat{A}||_{2}=\max_{|\psi\rangle\neq 0}\,||\hat{A}|\psi\rangle||_{2}/||\psi||_{2}$. The same numbers $n_{l}$ and $\xi_{l}$ apply equally to all quasi-degenerate eigenvalues in the perturbed system, hence, are independent of the index $l^{\prime}$ distinguishing these eigenvalues.
@@
The preceding two paragraphs outline the issues that we tackle in the present work. To resolve them, we will adopt the perspective of spectral response theory, and show, in the next section, how this leads to a natural generalization of the Petermann factor that reliably characterizes the mathematical nature and physical role of mode nonorthogonality both at and near exceptional points, as well as the connection between these two situations.
@@ III.1 Setting the scene
[…]
@@
[…]
@@ III.2 Petermann factor of exceptional points
We now connect this behavior to the mode nonorthogonality at the EP. For this, we first note that for rank-1 matrices such as $N^{n_{l}-1}_{l}$, the spectral norm gives the same result as the Frobenius norm $||A||_{\text{F}}=(\mbox{tr}\,A^{\dagger}A)^{1/2}$. Hence,
@@
$\hat{P}_{l}^{(R)}$ $=\sum_{km}A_{km}|J^{(l)}_{k}\rangle\langle J^{(l)}_{m}|,$ (36)
$\hat{P}_{l}^{(L)}$ $=\sum_{km}\tilde{A}_{km}|\tilde{J}^{(l)}_{k}\rangle\langle\tilde{J}^{(l)}_{m}|,$ (37)
@@
We therefore conclude that just as for isolated resonances, the physical response of systems at an EP is enhanced by mode nonorthogonality, which can be quantified by a pair of generalized Petermann factors $K_{l}^{(R,L)}$ that are directly related to the spectral response strength. The geometric content of this enhancement is emphasized by its explicit form (40), (41), while Eq. (39) affords the stated physical interpretation. Finally, these EP Petermann factors can be naturally combined into an overall EP Petermann factor $K_{l}$, expressed physically in terms of enhanced response in Eq. (42) and geometrically in terms of mode-nonorthogonality in Eq. (43).
@@ III.3 Consistency with the behavior near exceptional points
The relation between these projectors concerns both their structure, magnitude and their complex phase. For the structure, we observe that all projectors become aligned according to $\hat{P^{\prime}}_{l^{\prime}}\propto\hat{N}_{l}^{n_{l}-1}$, cf. Eqs. (10) and (44). As emphasized before, the eigenvectors in both expressions are subject to different biorthogonality conditions [see Eq. (13) for the Jordan chain]. For the magnitude of the ensuing proportionality constant, we note that according to Eq. (24), all eigenvalues have asymptotically identical Petermann factors $K_{l^{\prime}}=\mbox{tr}\,\hat{P^{\prime}}_{l^{\prime}}^{\dagger}\hat{P^{\prime}}_{l^{\prime}}\approx|\xi_{l}|^{2}/n_{l}^{2}|\Delta|^{2n_{l}-2}$. To furthermore capture the complex phase relations between all terms, we introduce the residues
@@
Utilizing the identity (see Appendix B)

### v2

@@ Abstract
The nonorthogonality of modes in open systems significantly modifies their resonant response, resulting in quantitative and qualitative deviations from Breit-Wigner resonance relations. For isolated resonances with a Lorentzian lineshape, the deviations amount to a quantitative enhancement of the resonance linewidth, given by the Petermann factor, which is determined by the overlap of left and right eigenmodes of the underlying effectively non-Hermitian Hamiltonian. The Petermann factor diverges at exceptional points (EPs), where complex resonance frequencies become degenerate, and right and left eigenmodes are orthogonal to each other. This divergence signifies a qualitative departure from a Lorentzian lineshape, which has moved into the focus of recent attention. In this work, we develop the analog of the Petermann factor for this qualitatively modified response at an EP, and describe how this EP Petermann factor manifests in a variety of physical scenarios. First, we identify this analog in physical terms as an enhancement of the response of a system to external or parametric perturbations. Utilizing two natural orthogonally projected reference systems based on the right and left eigenvectors, we show that each choice carries a precise geometric interpretation that naturally extends the notion of the Petermann factor for isolated resonances to EPs. The two choices can be combined into an overall EP Petermann factor, which again can be expressed in purely geometric terms. Second, we illuminate the geometric mechanisms that determine the size of the EP Petermann factor, by considering the role of modes participating in the degeneracy and those that remain spectrally separated from it. This also leads to a systematic description to evaluate this factor. Third, we design a system that allows us to study the EP Petermann factor in a specific physical setup, consisting of two microrings coupled to a waveguide with embedded semitransparent mirrors. With this example, we demonstrate that our approach gives more accurate results for the spectral response strength than conventional truncation schemes. These results complete the description of systems operating at exceptional points in the same way as the original Petermann factor does for isolated resonances, and pave the way to the design of open systems with unconventional spectral response, be it in emission or in response to static and dynamic perturbations.
@@ I Introduction
The field of non-Hermitian physics has attracted a considerable amount of attention in various subdisciplines, such as photonics [1, 2], acoustics [3], and condensed-matter physics [4]. The term “non-Hermitian” refers to the (effective) Hamiltonian [5, 6], which describes not only conservative but also dissipative dynamics of waves, stemming, e.g., from material gain, material loss, and radiation loss.
@@
The Petermann factor diverges to infinity [24, 25, 26] at so-called exceptional points (EPs) in parameter space [27]. At such an EP of order $n$, exactly $n$ eigenstates and their corresponding eigenvalues of the Hamiltonian coalesce. EPs are highly sensitive to perturbations: A perturbation of strength $\varepsilon>0$ results generically in eigenvalue changes proportional to the $n$th root of $\varepsilon$ [27]. This sensitivity makes EPs good candidates for sensing applications [28, 29, 30, 31, 32, 33, 34, 35]. The response of the system at the EP in terms of eigenvalue changes can be quantified by the spectral response strength $\xi$ [36].
The aim of this paper is to extend the concept of the Petermann factor, which was originally introduced for isolated resonances, to the degenerate situation at an EP. Physically, this EP Petermann factor quantifies the enhancement of the response strength above that of reference systems from which the mode-nonorthogonality is removed. For these systems, there are two natural choices: One pertains to the excitation of the system, while the other relates to the detection of the resulting response. The corresponding enhancement factors carry a precise geometric interpretation that naturally extends the notion of the Petermann factor for isolated resonances to EPs, and link the physical enhancement of the response strength to purely geometric information quantifying nonorthogonality of the right and left generalized eigenvectors. This information can further be combined into an overall EP Petermann factor, which again carries a precise geometric interpretation quantifying mode non-orthogonality. To substantiate the physical interpretation of these concepts, we illuminate the role of the EP Petermann factor in various contexts, such as system response to perturbations and quantum noise. In addition to these theoretical insights, we present a detailed prescription for determining the EP Petermann factor from both analytical and numerical models. Finally, we demonstrate the practical importance of the EP Petermann factor by comparing its predictions for the spectral response strength to those of standard truncation schemes. Through both an analytical toy model and a full numerical simulation of a realistic photonic system, we reveal how the inclusion of the EP Petermann factor leads to significantly more accurate results.
@@ II.1 Definitions and preliminary considerations
[…]
@@ II.2 Spectral decomposition of Hermitian matrices
We recall that for a Hermitian Hamiltonian, the eigenvalues $E_{l}$ are real, while the right and left eigenvector can be identified via $\langle L_{l}|=|R_{l}\rangle^{\dagger}$. Additionally, the eigenvectors of distinct eigenvalues $E_{k}\neq E_{l}$ are mutually orthogonal, $\langle R_{k}|R_{l}\rangle=0$, and individual eigenvectors can be normalized to obey $\langle R_{l}|R_{l}\rangle=1$. Furthermore, this construction can be straightforwardly extended to the case of degenerate eigenvalues. The mathematical reason is that in Hermitian systems, the algebraic multiplicity $n_{l}$, given by the multiplicity of roots in the secular equation (2), is identical to the geometric multiplicity $g_{l}$, given by the number of linearly independent eigenvectors admitted by Eq. (1). Employing a suitable orthogonal basis in the subspace spanned by these vectors, we therefore can construct an orthonormal basis for the complete system, with
$\langle R_{k}|R_{l}\rangle=\delta_{kl}\quad\mbox{(orthonormal basis)}.$ (3)
@@
$\hat{H}=\sum_{l}E_{l}|R_{l}\rangle\langle R_{l}|\quad\mbox{(Hermitian matrix)},$ (4)
@@
$\hat{P}_{l}=\sum_{m|E_{m}=E_{l}}|R_{m}\rangle\langle R_{m}|\quad\mbox{(Hermitian matrix)},$ (5)
@@
The spectra decomposition then takes the compact form
$\hat{H}=\sum_{l}E_{l}\hat{P}_{l}\quad\mbox{(Hermitian matrix)},$ (7)
@@ II.4 Exceptional points
Degeneracies of non-Hermitian systems differ from degeneracies in Hermitian systems in that along with the eigenvalues, the eigenvectors of the participating eigenvalues can also coalesce. For typical degeneracies in non-Hermitian systems, the geometric multiplicity is indeed smaller than the algebraic multiplicity, and in the generic case (i.e., without further fine tuning of parameters), the geometric multiplicity is one. These generic non-Hermitian degeneracies are called EPs. To cover these cases, we can view the indices $l$ as labels of separate eigenvectors, each associated with $n_{l}$ degenerate eigenvalues. This includes the case that the spectrum displays several EPs, while scenarios with higher geometric multiplicity can, in the context of the present work, be considered as additional accidental degeneracies between different EPs.
As the number of independent eigenvectors is reduced, these vectors do not form a complete basis (the Hamiltonian is then said to be defective). An important consequence is that the Hamiltonian can no longer be diagonalized. Instead, a complete basis can be obtained by supplementing each right eigenvector $|R_{l}\rangle$ by a Jordan chain of $n_{l}-1$ covectors $|J^{(l)}_{n}\rangle$, fulfilling
@@
[…]
@@
This biorthogonality condition will be used at EPs, where it replaces the conventional biorthogonality condition in Eq. (8). This universal rule applies throughout the remainder of the paper. The transformation into this biorthogonal basis corresponds to a Jordan decomposition
@@
where $D$ is the matrix with unit entries on the first superdiagonal, i.e., $D_{nn^{\prime}}=\delta_{n,n^{\prime}-1}$. An important feature of these matrices is that they are nilpotent with index $n_{l}$, hence, $D^{n_{l}}=0$ but $D^{n_{l}-1}\neq 0$. For an isolated eigenvalue, we simply have $D=0$, so that this term can be dropped from the decomposition. If this applies to all eigenvalues, this aligns with Eq. (10) after applying the similarity transformation (14).
@@
While the construction of the biorthogonal Jordan chains allows for some freedom, which is also reflected in the detailed form of the transformation matrix $T$, the operators $\hat{P}_{l}$ and $\hat{N}_{l}$ are unique. From their construction, the operators $\hat{P}_{l}$ are oblique projectors onto the (generalized) eigenspace associated with the specified eigenvalue $E_{l}$ that continue to fulfill Eq. (6), but now are of finite rank $n_{l}$. Furthermore, the operators $\hat{N}_{l}$ are nilpotent with index $n_{l}$, i.e., $\hat{N}_{l}^{n_{l}}=0$ but $\hat{N}_{l}^{n_{l}-1}\neq 0$. Additionally, these operators obey the important relation
@@ II.5 Measures of mode nonorthogonality for isolated resonances
The spectral decompositions described above imply a strong link between the spectral features of non-Hermitian systems and the geometric structure of the states associated with them. To prepare the discussion of these features at EPs, we survey common theoretical mode-nonorthogonality measures that illuminate different aspects of this link for spectrally isolated resonances, and detail how they are closely related to each other. The detailed physical contexts for these quantities are provided in later sections, where we generalize these notions to apply to systems at EPs.
@@
$K_{l}=\frac{(\sum_{m}|\langle m|R_{l}\rangle|^{2})^{2}}{|\sum_{m}\langle m|R_{l}\rangle^{2}|^{2}}\quad\mbox{(reciprocal systems).}$ (22)
@@
$r_{l}=\frac{|\sum_{m}\langle m|R_{l}\rangle^{2}|}{\sum_{m}|\langle m|R_{l}\rangle|^{2}}\quad\mbox{(reciprocal systems),}$ (23)
@@
Even though formulated from different perspectives, the different mode-nonorthogonality measures are therefore strictly related to each other. Hence, for isolated resonances, physical considerations of mode crosstalk, wavefunction imbalances, and eigenvalue sensitivity all relate to a very specific and universal aspects of mode nonorthogonality. In the following sections, we will explore how a richer and more complex picture emerges at EPs.
@@ II.6 Mode nonorthogonality near exceptional points
[…]
@@
The preceding two paragraphs outline the issues that we tackle in the present work. To resolve them, we will adopt the perspective of spectral response theory, and show, in the next section, how this leads to a natural generalization of the Petermann factor that reliably characterizes the mathematical nature and physical role of mode nonorthogonality both at and near EPs, as well as the connection between these two situations.
@@ III.1 Setting the scene
$K_{l}=\mbox{tr}\,\hat{P}_{l}^{\dagger}\hat{P}_{l},\quad\mbox{(isolated eigenvalue)}$ (26)
@@
$\mathcal{P}(E)\sim\frac{K_{l}}{|E-E_{l}|^{2}},\quad\mbox{(isolated eigenvalue)}$ (29)
@@ III.2 Petermann factor of exceptional points
We now connect this behavior to the mode nonorthogonality at the EP. For this, we first note that for rank-1 matrices such as $N^{n_{l}-1}_{l}$, the spectral norm gives the same result as the Frobenius norm $A||_{\text{F}}=(\mbox{tr}\,A^{\dagger}A)^{1/2}$. Hence,
@@
[…]
@@
We therefore conclude that just as for isolated resonances, the physical response of systems at an EP is enhanced by mode nonorthogonality,
$\mathcal{P}(E)\sim\frac{\left(\xi_{l}^{(R,L)}\right)^{2}K_{l}^{(R,L)}}{|E-E_{l}|^{2n_{l}}},\quad\mbox{(EP)}$ (48)
which can be quantified by a pair of generalized Petermann factors $K_{l}^{(R,L)}$ that are directly related to the spectral response strength. Concretely, these Petermann factors can then be isolated by comparison of total response power in Eq. (48) with the conditioned response powers
$\mathcal{P}_{R}(E)=\mathrm{tr}\,\hat{G}^{\dagger}(E)\hat{G}(E)\hat{P}^{(R)}\sim\frac{\left(\xi_{l}^{(R)}\right)^{2}}{|E-E_{l}|^{2n_{l}}},$ (49)
$\mathcal{P}_{L}(E)=\mathrm{tr}\,\hat{G}^{\dagger}(E)\hat{P}^{(L)}\hat{G}(E)\sim\frac{\left(\xi_{l}^{(L)}\right)^{2}}{|E-E_{l}|^{2n_{l}}},$ (50)
corresponding to a source covering the right eigenspace and a detector covering the left eigenspace, respectively. The geometric content of this enhancement is emphasized by its explicit form (40), (41), while Eq. (39) affords the stated physical interpretation. Finally, these EP Petermann factors can be naturally combined into an overall EP Petermann factor $K_{l}$, expressed physically in terms of enhanced response in Eq. (42) and geometrically in terms of mode-nonorthogonality in Eq. (43).
@@ III.3 Consistency with the behavior near exceptional points
The relation between these projectors concerns their structure, magnitude and their complex phase. For the structure, we observe that all projectors become aligned according to $\hat{P^{\prime}}_{l^{\prime}}\propto\hat{N}_{l}^{n_{l}-1}$, cf. Eqs. (10) and (44). As emphasized before, the eigenvectors in both expressions are subject to different biorthogonality conditions [see Eq. (13) for the Jordan chain]. For the magnitude of the ensuing proportionality constant, we note that according to Eq. (24), all eigenvalues have asymptotically identical Petermann factors $K_{l^{\prime}}=\mbox{tr}\,\hat{P^{\prime}}_{l^{\prime}}^{\dagger}\hat{P^{\prime}}_{l^{\prime}}\approx|\xi_{l}|^{2}/n_{l}^{2}|\Delta|^{2n_{l}-2}$. To furthermore capture the complex phase relations between all terms, we introduce the residues
@@
Utilizing the identity for an arbitrary complex constant $a$ (see Appendix B)